vector has y component ay= 13.0m

1
Expert's answer
2021-11-08T20:52:26-0500

Vector has y-component "a_y= +13.0m" . "a" makes an angle of "36.0\degree" counterclockwise from the "+y" -axis. (a) What is the "x" -component of "a"?


"\tan (90\degree+36\degree)=\dfrac{a_y}{a_x}"

"a_x=\dfrac{a_y}{\tan 126\degree}=-13\tan 36\degree m\approx-9.445m"

(b) What is the magnitude of "a"?


"|a|=\sqrt{a_x^2+a_y^2}=\sqrt{(-13\tan 36\degree m)^2+(13m)^2}"

"=\dfrac{13}{\cos 36\degree}m\approx16.069m"